OpenStudy (anonymous):

Tickets for a show cost 3 or 5 dollars. If 50 tickets were sold for a total of 230 how many tickets were sold for 3?

OpenStudy (anonymous):

there shud be something else given

OpenStudy (anonymous):

to find your answer solve \[5x+3(50-x)=230\]

OpenStudy (anonymous):

There's answer's to it to. 10,20,25,30,40 It's a multiple choice

OpenStudy (anonymous):

40

OpenStudy (anonymous):

How'd you get it though. ?

OpenStudy (anonymous):

5x - 150 - 3x = 230 2x = 80 x = 40

OpenStudy (anonymous):

*shud be +150

OpenStudy (anonymous):

Thankss!

OpenStudy (anonymous):

actually my equation was wrong since i am solving for $5 tickets. so be careful here

OpenStudy (anonymous):

it asks for number of $3 tickets sold

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so the answer is 40 $5 tickets and therefore 10 $3 tickets. the correct answer to the question asked is 10, not 40

OpenStudy (anonymous):

How'd you get 10?

OpenStudy (anonymous):

you could either solve \[3x+5(50-x)=230\] or reason that if there are 40 $5 tickets sold and the total is 50 then there must be 10 sold at $3

OpenStudy (anonymous):

Yeah. okay thanks!